A telemetry system is used to quantify kinematic
values of a ski jumper immediately before she
leaves the ramp. According to the system r =
500ft, ̇r = −105ft/sec, ̈r = −10 ft/sec2
, θ = 25◦
,

̇θ = 0.07rad/sec, ̈θ = 0.06rad/sec2

. Determine
(a) the velocity of the skier immediately before
she leaves the jump, (b) the acceleration of the
skier at this instant, (c) the distance of the jump
d neglecting lift and air resistance.
Answer #1


velocity of the skier V = Vrey + Voco () 20 V= rer + reco 25° Vro velocity of skier in 25° radial direction Vo: of Skier in vV= (-105 cos 25 i – las sin 25 j) + (-35 cos 65 i + 35 sin 65j) (-105 cos 25 - 35 cos 65) i + (-los sin 25 +35 Sin 65) (-109.* Rewriting the components rectangular by resolving a: acceleration ego in terms of recto polar components (-12-45 cos 25 i 1jo = 12.965 ft/s 9 = a + got + / t* 9: o + J2: 65 t +. - (32-2) + 9= 12.65 t + 16.1t2 * Cos 30° d X = dcos 30° Sin 30 - y-loconsidering positive value for t, 109.95 (3.343) 367. 569 Ft X from eq^ a & d COS 30° 367. 569- d cos 30 d=424.43 ft CS Scan